1. Represent the function 10/1−10x as a power series f(x)=∞∑n=0cn x^n
Compute the first few coefficients of this power series:
c0=   
c1=   
c2=   
c3=   
c4=
Find the radius of convergence R=

2. The Taylor series for f(x)=e^x at a = 2 is ∞∑n=0 cn(x−2)^n.
Find the first few coefficients.
c0=   
c1=   
c2=
c3=
c4=
Answer #1

@ We serits for know the I 1-x no power an Š (10x) (qla xay for} I-102 no Ž 10n.am I-lox No multiply both sides by io & 8=ea dx x=2 We have, f(x) = ex f (x) = f(x) = f(x) = f(x) = en {: d (ex Thus, f (2) = f(2) = f (2) = f(al = f (2) - e² pow